OpenStudy (anonymous):

I just need help with this problem. I just don't know where to begin. Given the system: 2x-3y+4z=21 5x+6y+8z=10 Z=x+y What is the value of x?

OpenStudy (anonymous):

move the x and y over to the left side on the bottom equation giving z - x - y = 0

OpenStudy (anonymous):

or in a better order -x - y + z = 0

OpenStudy (anonymous):

What should I do from there.

OpenStudy (anonymous):

lets call each equation R1, R2, R3 so we have 2x-3y+4z=21 R1 5x+6y+8z=10 R2 - x - y + z = 0 R3 we need to get to a situation where we have an equation with one variable

OpenStudy (anonymous):

multiply R1 by 5/2, and then subtract R1 - R2 = R2a so you obtain a new formula R2a with only two variables y and z.

OpenStudy (anonymous):

then you need to multiply R1 by somthing so you can subtract R1 - R3 = R3a R3a being a new formula again with only y and z.

OpenStudy (anonymous):

then you multiply R3a by something in order that you can cancel the y's when you do R3a - R2a = R3b you should be left with an equation with one variable which you can solve

OpenStudy (anonymous):

and then plug that value back into R3a or R2a which have two variables, solve that. And then plug both the y value and z into R1 to find x

OpenStudy (anonymous):

This process is called gaussian elimination

OpenStudy (anonymous):

Alright bin working in a similar question like this for my exam review and I think I just got stumped but If I have any more questions ill post on this

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Than you for your help

OpenStudy (anonymous):

np